OpenStudy (lori):

Problem: At the bakery each scone cost $0.95 and each muffin costs $0.80. If Alex needs to buy 30 pastries for his class(a combination of both), how many of each will he buy if he has $30.00 to spend?

OpenStudy (anonymous):

Hi there. Are you interested in learing the material or just looking for an answer to copy?

OpenStudy (lori):

learning I have a college math final tuesday

OpenStudy (anonymous):

0.95s + 0.8m = 30 we know he'll buy more than thirty since they cost less than 1 dollar each he will save 5 cents per scone and 20 cents per muffin

OpenStudy (anonymous):

OK, this is a system of linear equations. Do you know how to solve a system of linear equations?

OpenStudy (lori):

yes

OpenStudy (anonymous):

Great; so, we just have to set up the system. Number of scones is s. Number of muffins is m. So we know first: Total pastries = 30, so s + m = 30 Total cost is $30, so 0.95 * s + 0.80 * m = $30 Now just simultaneously solve those two.

OpenStudy (lori):

solve with substitution?

OpenStudy (anonymous):

You can solve any way you want; they'll all give you the same answer. But, yeah, substitution is probably easiest for this particular problem.

OpenStudy (anonymous):

e.g. isolate s in equation one, then substitute that in where s appears in Eq. 2. You'll end up with an equation that has only one variable, m.

OpenStudy (lori):

I got m = 40 but it's not making sense to me

OpenStudy (anonymous):

so basically, after buying 15 muffins he will have gained 3 dollars b/c they were only 80 cents and he will have saved 75 cents from the 15 scones w/ 3.75 he can buy 4 muffins w/ a few pennies left over or 3 scones w/ a few pennies

OpenStudy (anonymous):

Haha, I just worked the problem and I got the same. It's broken. Are you sure the total cost is supposed to be $30?

OpenStudy (anonymous):

total he can buy max abt 34 if he wants to get a fairly equal amount of pastries

OpenStudy (anonymous):

make sure you gave us all the right numbers in the problem.

OpenStudy (lori):

I did it's on a practice final from my professor and it just confused me

OpenStudy (anonymous):

Since scones and muffins both cost less than one dollar, there's no combination of 30 muffins and scones that costs $30. It'll always cost less.

OpenStudy (anonymous):

look at what I did, I basically solved both as if they were a dollar and then found what I would have saved if they cost the correct amount

OpenStudy (lori):

I will bob06

OpenStudy (lori):

the second half of the question is as follows: use the previous part to model the problem using an inequality. List the possible number of scones that Alex could have purchased according to your model.....ugh!

OpenStudy (anonymous):

the problem is broken unless we use a larger number of pastries or a smaller number of dollars to spend

OpenStudy (lori):

could you guide me on how to do the second half of the question